please help me with similar triangles
ABC and DEF are similar What is the remaining side of a=10, d=12,c=6,d=6

Answer by venugopalramana(3286)

ABC and DEF are similar What is the remaining side of a=10, d=12,c=6,d=6
there is a mistake in data ..it is taken as a=10, b=12,c=6,d=6..you can check the answer and correct any way later.
in similar triangles we have
1.corresponding angles are equal and
2.corresponding sides are in proportion.
we are given that in triangle ABC...a=10,b=12,c=6,and
we are given that in triangle DEF...d=6,e=??,f=??
taking corresponding sides are in proportion,we get a/d=b/e=c/f
10/6=12/e=6/f
10/6=12/e or 10e=12*6=72 or e=72/10
10/6=6/f or 10f=6*6=36 or f=36/10